Testing:

H0:p=0.42H0:p=0.42
H1:p>0.42H1:p>0.42

Your sample consists of 143 subjects, with 62 successes. Calculate the test statistic, rounded to 2 decimal places

Hint: Round your ˆpp^ calculation to at least 4 decimal places before calculating the test statistic.

z=z=
Answer #1

solution

this is right tailed test

\hat p = x / n = 62/143=0.4336

P0=0.42

1 - P0 = 1-0.42

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.4336 -0.42/ [\sqrt(0.42*0.58) /143 ]

Z=0.33